OpenStudy (anonymous):

What is the answer: 4x2 - 16=8 ?

OpenStudy (chaise):

\[4x ^{2}-16=8\] 4x^2 = 24 x^2 = 6 x = sqrt(6) x = +/- 2.5

OpenStudy (akshay_budhkar):

silly mistake lol

OpenStudy (chaise):

Haha, I thought so, it's fine. we all make them :)

OpenStudy (akshay_budhkar):

x = sqrt(6) or x= - sqrt 6